### v2

@@ 8 Empirical Application: Finite Horizon DDC
This section applies our framework to a finite-horizon dynamic labor supply model for taxi drivers in New York City (NYC). In the model, the market-level supply shock is the latent variable. Our reference distribution is an AR(1) process. We consider the sensitivity analysis of the labor supply elasticity with respect to such distributional assumptions.
@@ 8.5 Sensitivity Analysis
$\sup_{g^{2}_{tk}}\mathbb{E}_{F}\mathbb{E}_{w_{t+1},w_{t}}\left[\psi_{t}(k,\xi_{t},\xi_{t+1},w_{t},w_{t+1};u^{\prime},p_{t},p_{t+1},g^{2}_{tk})\right]\text{ for }\forall\ t\leq T-1,k$
@@ 9 Conclusion
We propose a computationally tractable framework to quantify the sensitivity of outcomes of interest to distributional assumptions about the serial dependence of latent variables in dynamic structural models. We derive bounds on a scalar parameter of interest by perturbing a reference distribution, while imposing a stationarity condition for time-homogeneous models or a Markovian condition for time-inhomogeneous models. We apply the approach to an infinite-horizon dynamic demand model for new cars in the UK, Germany, and France, and a finite-horizon dynamic labor supply model for taxi drivers in New York City.
@@ Appendix A Additional Examples
The model-implied conditional choice probability is given by:
$p(j|x,I,\xi)=\frac{\exp\left(v_{j}(x,I,\xi)\right)}{\sum_{j^{\prime}\in\mathcal{J}}\exp\left(v_{j^{\prime}}(x,I,\xi)\right)}$
@@
We rewrite the structural constraints as:
where $\psi$ is the sum of the two expressions inside the expectations and $g:=(g^{1},g^{2})$.
@@
where $P_{0}(j|x,I)$ is the population CCP and $q_{0}(j,x,I)$ is the population continuous choice function. We discretize $\mathcal{X}\times\mathcal{I}$ for estimation, and rewrite the moment conditions as:
$\mathbb{E}_{F}\left[m(U;v)\right]=P_{0}$
where $m$ ($P_{0}$) stacks the model-implied (population) CCPs and continuous choice functions.

### v3

@@ 8 Empirical Application: Finite Horizon DDC
This section applies our framework to a finite-horizon dynamic labor supply model for taxi drivers in New York City (NYC). In the model, the market-level supply shock is the latent variable. Our reference dynamic process is an AR(1) process. We consider the sensitivity analysis of the labor supply elasticity with respect to such distributional assumption.
@@ 8.5 Sensitivity Analysis
$\sup_{g^{2}_{tk}}\mathbb{E}_{F}\mathbb{E}_{w_{t+1},w_{t}}\left[\psi_{t}(k,\xi_{t},\xi_{t+1},w_{t},w_{t+1};u^{\prime},p_{t}^{\prime},p_{t+1}^{\prime},g^{2}_{tk})\right]\text{ for }\forall\ t\leq T-1,k$
@@ 9 Conclusion
We propose a computationally tractable framework to quantify the sensitivity of outcomes of interest to misspecified latent-state dynamics in structural models. We derive bounds on a scalar parameter of interest by perturbing a reference dynamic process, while imposing a stationarity condition for time-homogeneous models or a Markovian condition for time-inhomogeneous models. We apply the approach to an infinite-horizon dynamic demand model for new cars in the UK, Germany, and France, and a finite-horizon dynamic labor supply model for taxi drivers in New York City.
@@ Appendix A Additional Examples
The model-implied conditional choice probability is $p(j|x,I,\xi)=\frac{\exp\left(v_{j}(x,I,\xi)\right)}{\sum_{j^{\prime}\in\mathcal{J}}\exp\left(v_{j^{\prime}}(x,I,\xi)\right)}$.
@@
We rewrite the structural constraints as $\sup_{g\in\mathcal{G}}\mathbb{E}_{F}\left[\psi(U;\theta,v,g)\right]=0$ where $\psi$ is the sum of the two expressions inside the expectations and $g:=(g^{1},g^{2})$.
@@
where $P_{0}(j|x,I)$ is the population CCP and $q_{0}(j,x,I)$ is the population continuous choice function. We discretize $\mathcal{X}\times\mathcal{I}$ for estimation, and rewrite the moment conditions as $\mathbb{E}_{F}\left[m(U;v)\right]=P_{0}\mathbb{E}_{F}\left[m(U;v)\right]=P_{0}$ where $m$ ($P_{0}$) stacks the model-implied (population) CCPs and continuous choice functions.